# Distributionally Robust Policy Learning under Concept Drifts

### v1

@@ Abstract
Distributionally robust policy learning aims to find a policy that performs well under the worst-case distributional shift, and yet most existing methods for robust policy learning consider the worst-case joint distribution of the covariate and the outcome. The joint-modeling strategy can be unnecessarily conservative when we have more information on the source of distributional shifts. This paper studies a more nuanced problem — robust policy learning under the concept drift, when only the conditional relationship between the outcome and the covariate changes. To this end, we first provide a doubly-robust estimator for evaluating the worst-case average reward of a given policy under a set of perturbed conditional distributions. We show that the policy value estimator enjoys asymptotic normality even if the nuisance parameters are estimated with a slower-than-root-$n$ rate. We then propose a learning algorithm that outputs the policy maximizing the estimated policy value within a given policy class $\Pi$, and show that the sub-optimality gap of the proposed algorithm is of the order $\kappa(\Pi)n^{-1/2}$, with $\kappa(\Pi)$ is the entropy integral of $\Pi$ under the Hamming distance and $n$ is the sample size. A matching lower bound is provided to show the optimality of the rate. The proposed methods are implemented and evaluated in numerical studies, demonstrating substantial improvement compared with existing benchmarks.
@@ 1 Introduction
In a wide range of fields, the abundance of user-specific historical data provides opportunities for learning efficient individualized policies. Examples include learning the optimal personalized treatment from electronic health record data (Murphy,, 2003; Kim et al.,, 2011; Chan et al.,, 2012), or obtaining an individualized advertising strategy using past customer behavior data (Bottou et al.,, 2013; Kallus and Udell,, 2016). Driven by such a practical need, a line of works have been devoted to developing efficient policy learning algorithms using historical data — a task often known as offline policy learning (Dudík et al.,, 2011; Zhang et al.,, 2012; Swaminathan and Joachims, 2015a, ; Swaminathan and Joachims, 2015b, ; Swaminathan and Joachims, 2015c, ; Kitagawa and Tetenov,, 2018; Athey and Wager,, 2021; Zhou et al.,, 2023; Zhan et al.,, 2023; Bibaut et al.,, 2021; Jin et al.,, 2021; Jin et al., 2022a, ).
@@
Most existing methods for offline policy learning deliver performance guarantees under the premise that the target environment remains the same as that from which the historical data is collected. It has been widely observed, however, that such a condition is hardly met in practice (see e.g., Recht et al., (2019); Namkoong et al., (2023); Liu et al., (2023); Jin et al., (2023) and the references therein). Under distribution shift, a policy learned in one environment often shows degraded performance when deployed in another environment. To address this issue, there is an emerging body of research on robust policy learning, which aims at finding a policy that still performs well when the target distribution is perturbed. Pioneering works in this area consider the case where the joint distribution of the covariates and the outcome is shifted from the training distribution, and researchers devise algorithms that output a policy achieving reliable worst-case performance under the aforementioned shifts (Si et al.,, 2023; Kallus et al.,, 2022). The joint modeling approach, however, ignores the type of distributional shifts, and the resulting worst-case value can be unnecessarily conservative in practice.
@@
Indeed, distributional shifts can be categorized into two classes by their sources: (1) the shift in the covariate $X$, and/or (2) the shift in the conditional relationship between the outcome $Y$ and the covariate $X$. The two types of distributional shifts are different in nature, have different implications on the objectives, and call for distinct treatment (Namkoong et al.,, 2023; Liu et al.,, 2023; Jin et al.,, 2023; Ai and Ren,, 2024). To be concrete, imagine that the distribution of covariates changes while that of $Y\,|\,X$ remains invariant — in this case, the distribution shift is identifiable/estimable since the covariates are often accessible in the target environment. As a result, it is often unnecessary to account for the worst-case covariate shift rather than directly correcting for it. Alternatively, when the $Y\,|\,X$ distribution changes but the $X$ distribution remains invariant, the distribution shift is no longer identifiable, where we can instead apply the worst-case consideration to guarantee performance. This latter setting, known as concept drift, occurs due to sudden external shocks (Widmer and Kubat,, 1996; Lu et al.,, 2018; Gama et al.,, 2014). For example, in advertising, the customer behavior can evolve over time as the environment changes, while the population remains largely the same. In personalized product recommendation, similar population segments in developed and emerging markets may prefer different product features. In these applications, with the one extra bit of information that the shift is only in the conditional reward distribution, can we obtain a more efficient policy learning algorithm?
@@
Motivated by the above situations, we study robust policy learning under concept drift in this work. Most existing methods for robust policy learning (Si et al.,, 2023; Kallus et al.,, 2022) model the distributional shift jointly without distinguishing the sources, and the corresponding algorithms turn out to be suboptimal. The reason behind their suboptimality is that the worst-case distributions under the two models — the joint-shift model and the concept-drift model — can be substantially different, so it would be a “waste” of our budget to consider adversarial distributions that are not feasible under concept drift. It is worth mentioning that a recent paper by Mu et al., (2022) accounts for the sources of distributional shifts in policy learning; their approach, however, applies only when the covariates take a finite number of values, and therefore is limited in its applicability. When the covariate space is infinite, it remains unclear how to efficiently learn a robust policy under concept drift. The current work aims to fill in the gap by answering the following question:
How can we efficiently learn a policy with optimal worst-case average performance under concept drift with minimal assumptions?
We provide a rigorous answer to the above question. Specifically, we assume the covariate distribution remains the same in the training and target environments, [^1] Otherwise, the covariate shift can be easily adjusted by covariate matching discussed earlier. while the $Y\,|\,X$ distribution shift is bounded in KL-divergence by a pre-specified constant $\delta$. Our goal is to find a policy that maximizes the worst-case averaged outcome over all possible target distributions satisfying the previous condition.
@@ 1.1 Our contributions
  1. 1.
    Policy evaluation: Given a policy, we present a doubly-robust estimator for the worst-case policy value under concept drift. We prove that the estimator is asymptotic normal under mild conditions on the estimation rate of the nuisance parameter. Our approach involves solving the dual form of a distributionally robust optimization problem and taking a de-biased step to deal with the slow convergence of the optimizer, thereby obtaining an estimator with root-n convergence rate.
  2. 2.
    Policy learning: We propose a robust policy learning algorithm that outputs a policy maximizing the estimated policy value over a policy class $\Pi$. Compared with the oracle optimal policy, the policy provided by our algorithm with high probability has a regret/suboptimality gap of the order $\kappa(\Pi)/\sqrt{n}$, where $\kappa(\Pi)$ is a measure quantifying the policy class complexity (to be formalized shortly) and $n$ is the number of samples. Compared with Mu et al., (2022), our algorithm and theory apply to general covariate spaces and potentially infinite policy classes, while their method is restricted to finite covariate space and policy class. We complement the upper bound with a matching lower bound, thus establishing the minimax optimality of our proposed algorithm. We summarize the comparison between our result and prior work in Table 1 for better demonstration.
  3. 3.
    Implementation and empirics: We provide efficient implementation of our robust policy learning algorithm, and compare its empirical performance with existing benchmarks in numerical studies. Our proposed method exhibits substantial improvement.
Athey and Wager, (2021)
Zhou et al., (2023)
Si et al., (2023) Joint
Kallus et al., (2022) Joint
Mu et al., (2022) Separate $O\big{(}\sqrt{\frac{\log n\log(|\mathcal{X}||\mathcal{A}|)}{n}}\big{)}$
Table 1: Comparison of results in the offline policy learning literature. “Unknown $\pi_{0}$” refers to whether an algorithm assumes knowledge of the behavior policy $\pi_{0}$. “General $\mathcal{X}$” refers to whether an algorithm allows for general types of covariates. Athey and Wager, (2021); Zhou et al., (2023); Si et al., (2023); Kallus et al., (2022) have the regret upper and lower bounds for the specific problems they consider that are not directly comparable to ours, so we do not include them in the table. $|\mathcal{X}|$ refers to the cardinality of the covariate support (if finite) and $|\mathcal{A}|$ to that of the action set. $\kappa(\Pi)$ and $\text{Ndim}(\Pi)$ are the entropy integral under Hamming distance and the Natarajan dimension of a policy class $\Pi$, with the relation $\kappa(\Pi)=O(\log(d)\text{Ndim}(\Pi))$, where $d$ is the dimension of the covariate space.
@@ 1.2 Related works
Offline policy learning.
There is a long list of works devoted to offline policy learning. Most of them assume no distributional shifts (e.g., Dudík et al., (2011); Zhang et al., (2012); Swaminathan and Joachims, 2015a ; Swaminathan and Joachims, 2015b ; Swaminathan and Joachims, 2015c ; Kitagawa and Tetenov, (2018); Athey and Wager, (2021); Zhou et al., (2023)). Zhan et al., (2023); Jin et al., (2021); Jin et al., 2022a allow the data to be adaptively collected, but the distribution over the covariate and the (potential) outcomes remain invariant in the training and target environment.
As mentioned earlier, the work of Si et al., (2023); Kallus et al., (2022) study robust policy learning when the joint distribution of $(X,Y)$ ranges in the neighborhood of the training distribution; Mu et al., (2022) consider the case when the covariate shift and $Y\,|\,X$ shift are specified separately; their method, however, is restricted to finite covariate space, and their sub-optimality gap is logarithmic factors slower than parametric rates. More recently, Guo et al., (2024) considers a pure covariate shift with a focus on policy evaluation, where the setup and the goal are different from ours.
Distributionally robust optimization.
More broadly, our work is also closely related to DRO, where the goal is to learn a model that has good performance under the worst-case distribution (e.g., Bertsimas and Sim, (2004); Delage and Ye, (2010); Hu and Hong, (2013); Duchi et al., (2019); Dudík et al., (2011); Zhang et al., (2023)). The major focus of the aforementioned works involves parameter estimation and prediction in supervised settings; we however take a decision-making perspective and aim at learning an individualized policy with optimal worst-case performance guarantees.
1.3 Notation
@@
2 Preliminaries
Consider a set of $M$ actions denoted by $[M]$ and let $\mathcal{X}\subseteq\mathbb{R}^{d}$. Throughout the paper, we follow the potential outcome framework (Imbens and Rubin,, 2015), where $Y(a)\in\mathcal{Y}_{a}\subseteq\mathbb{R}$ denotes the potential outcome had action $a$ been taken for any $a\in[M]$. We posit the underlying data-generating distribution $P$ on the joint covariate-outcome random vector $(X,Y(1),\cdots,Y(M))\in\mathcal{X}\times\prod_{a=1}^{M}\mathcal{Y}_{a}$. Consider a data set $\mathcal{D}=\{(X_{i},A_{i},Y_{i})\}_{i\in[n]}$ consisting of $n$ i.i.d. draws of $(X,A,Y)$, where $X_{i}\in\mathcal{X}$ is the observed contextual vector, $A_{i}\in[M]$ the action, and $Y_{i}=Y(A_{i})$ the realized reward. The actions are selected by the behavior policy $\pi_{0}$, where $\pi_{0}(a\,|\,x):=\mathbb{P}(A_{i}=a\,|\,X=x)$ is the propensity score, for any $a\in[M],x\in\mathcal{X}$. We make the following assumptions for $\pi_{0}$ and $P$.
Assumption 2.1.
The behavior policy $\pi_{0}$ and the joint distribution $P$ satisfy the following.
  1. (1)
[…]
  2. (2)
    Overlap: for some $\varepsilon>0$, $\pi_{0}(a\,|\,x)\geq\varepsilon$, for all $(a,x)\in[M]\times\mathcal{X}$.
  3. (3)
    Bounded reward support: there exists $\bar{y}>0$, such that $0\leq Y(a)\leq\bar{y}$ for all $a\in[M]$.
The above assumptions are standard in the literature (see e.g., Athey and Wager,, 2021; Zhou et al.,, 2023; Si et al.,, 2023; Kallus et al.,, 2022). In particular, the unconfoundedness assumption guarantees identifiability, and the overlap assumption ensures sufficient exploration when collecting the training dataset. The bounded reward support is assumed for the ease of exposition, and can be relaxed to the sub-Gaussian reward straightforwardly.
@@ 2.1 The KL-distributionally robust formulation
We define an uncertainty set of neighboring distributions around $P$, whose conditional outcome distribution is bounded in KL divergence from $P$. Given a radius $\delta>0$, the uncertainty set of the conditional distribution is defined as $\mathcal{P}(P_{Y\,|\,X},\delta):=\big{\{}Q_{Y\,|\,X}:D_{\text{KL}}(Q_{Y\,|\,X}\,\|\,P_{Y\,|\,X})\leq\delta\big{\}}$, where $P_{Y\,|\,X}$ and $Q_{Y\,|\,X}$ refers to the distribution of $(Y(1),\dots,Y(d))\,|\,X$ under $P$ and $Q$ respectively. The distributionally robust policy value for any policy $\pi$ at level $\delta$ is defined as
@@
Under this formulation, our goal is to learn a “robust” policy with a high value of $\mathcal{V}_{\delta}(\pi)$ using a dataset drawn from $P$. The task here is two-fold: we need to (i) estimate the policy value $\mathcal{V}_{\delta}(\pi)$ for a given policy $\pi$, and (ii) find a near-optimal robust policy $\widehat{\pi}\in\Pi$ whose policy value is close to the optimal policy $\pi^{*}_{\delta}$. Here, the performance of a learned policy $\widehat{\pi}$ is measured by the sub-optimality gap (regret), defined as

### v2

@@ Abstract
Distributionally robust policy learning aims to find a policy that performs well under the worst-case distributional shift, and yet most existing methods for robust policy learning consider the worst-case joint distribution of the covariate and the outcome. The joint-modeling strategy can be unnecessarily conservative when we have more information on the source of distributional shifts. This paper studies a more nuanced problem — robust policy learning under the concept drift, when only the conditional relationship between the outcome and the covariate changes. To this end, we first provide a doubly-robust estimator for evaluating the worst-case average reward of a given policy under a set of perturbed conditional distributions. We show that the policy value estimator enjoys asymptotic normality even if the nuisance parameters are estimated with a slower-than-root-$n$ rate. We then propose a learning algorithm that outputs the policy maximizing the estimated policy value within a given policy class $\Pi$, and show that the sub-optimality gap of the proposed algorithm is of the order $\kappa(\Pi)n^{-1/2}$, where $\kappa(\Pi)$ is the entropy integral of $\Pi$ under the Hamming distance and $n$ is the sample size. A matching lower bound is provided to show the optimality of the rate. The proposed methods are implemented and evaluated in numerical studies, demonstrating substantial improvement compared with existing benchmarks.
Machine Learning, ICML
[…]
[Ruohan]rzredorange
Athey & Wager (2021)
Zhou et al. (2023)
Si et al. (2023) Joint
Kallus et al. (2022) Joint
Mu et al. (2022) Separate $O\big{(}\sqrt{\frac{\log n\log(|\mathcal{X}||\mathcal{A}|)}{n}}\big{)}$
Table 1: Comparison of results in the offline policy learning literature. “Unknown $\pi_{0}$” refers to whether an algorithm assumes knowledge of the behavior policy $\pi_{0}$. “General $\mathcal{X}$” refers to whether an algorithm allows for general types of covariates. Athey & Wager (2021); Zhou et al. (2023); Si et al. (2023); Kallus et al. (2022) have the regret upper and lower bounds for the specific problems they consider that are not directly comparable to ours, so we do not include them in the table. $|\mathcal{X}|$ refers to the cardinality of the covariate support (if finite) and $|\mathcal{A}|$ to that of the action set. $\kappa(\Pi)$ and $\text{Ndim}(\Pi)$ are the entropy integral under Hamming distance and the Natarajan dimension of a policy class $\Pi$, with the relation $\kappa(\Pi)=O(\log(d)\text{Ndim}(\Pi))$, where $d$ is the dimension of the covariate space.
@@ 1 Introduction
In a wide range of fields, the abundance of user-specific historical data provides opportunities for learning efficient individualized policies. Examples include learning the optimal personalized treatment from electronic health record data (Murphy, 2003; Kim et al., 2011; Chan et al., 2012), or obtaining an individualized advertising strategy using past customer behavior data (Bottou et al., 2013; Kallus & Udell, 2016). Driven by such a practical need, a line of works have been devoted to developing efficient policy learning algorithms using historical data — a task often known as offline policy learning (Dudík et al., 2011; Zhang et al., 2012; Swaminathan & Joachims, 2015a, b, c; Kitagawa & Tetenov, 2018; Athey & Wager, 2021; Zhou et al., 2023; Zhan et al., 2023; Bibaut et al., 2021; Jin et al., 2021, 2022a).
@@
Most existing methods for offline policy learning deliver performance guarantees under the premise that the target environment remains the same as that from which the historical data is collected. It has been widely observed, however, that such a condition is hardly met in practice (see e.g., Recht et al. (2019); Namkoong et al. (2023); Liu et al. (2023); Jin et al. (2023) and the references therein). Under distribution shift, a policy learned in one environment often shows degraded performance when deployed in another environment. To address this issue, there is an emerging body of research on robust policy learning, which aims at finding a policy that still performs well when the target distribution is perturbed. Pioneering works in this area consider the case where the joint distribution of the covariates and the outcome is shifted from the training distribution, and researchers devise algorithms that output a policy achieving reliable worst-case performance under the aforementioned shifts (Si et al., 2023; Kallus et al., 2022). The joint modeling approach, however, ignores the type of distributional shifts, and the resulting worst-case value can be unnecessarily conservative in practice.
@@
Indeed, distributional shifts can be categorized into two classes by their sources: (1) the shift in the covariate $X$, and/or (2) the shift in the conditional relationship between the outcome $Y$ and the covariate $X$. The two types of distributional shifts are different in nature, have different implications on the objectives, and call for distinct treatment (Namkoong et al., 2023; Liu et al., 2023; Jin et al., 2023; Ai & Ren, 2024). To be concrete, imagine that the distribution of covariates changes while that of $Y\,|\,X$ remains invariant — in this case, the distribution shift is identifiable/estimable since the covariates are often accessible in the target environment. As a result, it is often unnecessary to account for the worst-case covariate shift rather than directly correcting for it. Alternatively, when the $Y\,|\,X$ distribution changes but the $X$ distribution remains invariant, the distribution shift is no longer identifiable, where we can instead apply the worst-case consideration to guarantee performance. This latter setting, known as concept drift, occurs due to sudden external shocks (Widmer & Kubat, 1996; Lu et al., 2018; Gama et al., 2014). For example, in advertising, the customer behavior can evolve over time as the environment changes, while the population remains largely the same. In personalized product recommendation, similar population segments in developed and emerging markets may prefer different product features. In these applications, with the one extra bit of information that the shift is only in the conditional reward distribution, can we obtain a more efficient policy learning algorithm?
@@
Motivated by the above situations, we study robust policy learning under concept drift. Most existing methods for robust policy learning (Si et al., 2023; Kallus et al., 2022) model the distributional shift jointly without distinguishing the sources, and the corresponding algorithms turn out to be suboptimal. The reason behind their suboptimality is that the worst-case distributions under the two models — the joint-shift model and the concept-drift model — can be substantially different, so it would be a “waste” of our budget to consider adversarial distributions that are not feasible under concept drift. It is worth mentioning that a recent paper by Mu et al. (2022) accounts for the sources of distributional shifts in policy learning; their approach, however, applies only when the covariates take a finite number of values, and therefore is limited in its applicability. When the covariate space is infinite, it remains unclear how to efficiently learn a robust policy under concept drift. The current work aims to fill in the gap by answering the question:How can we efficiently learn a policy with optimal worst-case average performance under concept drift with minimal assumptions? We provide a rigorous answer to the above question. Specifically, we assume the covariate distribution remains the same in the training and target environments, [^1] Otherwise, the covariate shift can be easily adjusted by covariate matching discussed earlier. while the $Y\,|\,X$ distribution shift is bounded in KL-divergence by a pre-specified constant $\delta$. Our goal is to find a policy that maximizes the worst-case averaged outcome over all possible target distributions satisfying the previous condition.
@@ 1.1 Our Contributions
Policy Evaluation. Given a policy, we present a doubly-robust estimator for the worst-case policy value under concept drift. We prove that the estimator is asymptotic normal under mild conditions on the estimation rate of the nuisance parameter. Our approach involves solving the dual form of a distributionally robust optimization problem and taking a de-biased step to deal with the slow convergence of the optimizer, thereby obtaining an estimator with root-n convergence rate.
Policy Learning. We propose a robust policy learning algorithm that outputs a policy maximizing the estimated policy value over a policy class $\Pi$. Compared with the oracle optimal policy, the policy provided by our algorithm with high probability has a regret/suboptimality gap of the order $\kappa(\Pi)/\sqrt{n}$, where $\kappa(\Pi)$ is a measure quantifying the policy class complexity (to be formalized shortly) and $n$ is the number of samples. Compared with Mu et al. (2022), our algorithm and theory apply to general covariate spaces and potentially infinite policy classes, while their method is restricted to finite covariate space and policy class. We complement the upper bound with a matching lower bound, thus establishing the minimax optimality of our proposed algorithm. We summarize the comparison between our result and prior work in Table 1 for better demonstration.
Implementation and Empirics. We provide efficient implementation of our robust policy learning algorithm, and compare its empirical performance with existing benchmarks in numerical studies. Our proposed method exhibits substantial improvement.
@@ 1.2 Related Works
Offline Policy Learning. There is a long list of works devoted to offline policy learning. Most of them assume no distributional shifts (e.g., Dudík et al. (2011); Zhang et al. (2012); Swaminathan & Joachims (2015a, b, c); Kitagawa & Tetenov (2018); Athey & Wager (2021); Zhou et al. (2023)). Zhan et al. (2023); Jin et al. (2021, 2022a) allow the data to be adaptively collected, but the distribution over the covariate and the (potential) outcomes remain invariant in the training and target environment.
As mentioned earlier, Si et al. (2023); Kallus et al. (2022) study robust policy learning when the joint distribution of $(X,Y)$ ranges in the neighborhood of the training distribution; Mu et al. (2022) consider the case when the covariate shift and $Y\,|\,X$ shift are specified separately; their method, however, is restricted to finite covariate space, and their sub-optimality gap is logarithmic factors slower than parametric rates. More recently, Guo et al. (2024) considers a pure covariate shift with a focus on policy evaluation, where the setup and the goal are different from ours.
Distributionally Robust Optimization. More broadly, our work is closely related to DRO, where the goal is to learn a model that has good performance under the worst-case distribution (e.g., Bertsimas & Sim (2004); Delage & Ye (2010); Hu & Hong (2013); Duchi et al. (2019); Dudík et al. (2011); Zhang et al. (2023)). The major focus of the aforementioned works involves parameter estimation and prediction in supervised settings; we however take a decision-making perspective and aim at learning an individualized policy with optimal worst-case performance guarantees.
2 Preliminaries
Consider a set of $M$ actions denoted by $[M]$ and let $\mathcal{X}\subseteq\mathbb{R}^{d}$. Throughout the paper, we follow the potential outcome framework (Imbens & Rubin, 2015), where $Y(a)\in\mathcal{Y}_{a}\subseteq\mathbb{R}$ denotes the potential outcome had action $a$ been taken for any $a\in[M]$. We posit the underlying data-generating distribution $P$ on the joint covariate-outcome random vector $(X,Y(1),\cdots,Y(M))\in\mathcal{X}\times\prod_{a=1}^{M}\mathcal{Y}_{a}$. Consider a data set $\mathcal{D}=\{(X_{i},A_{i},Y_{i})\}_{i\in[n]}$ consisting of $n$ i.i.d. draws of $(X,A,Y)$, where $X_{i}\in\mathcal{X}$ is the observed contextual vector, $A_{i}\in[M]$ the action, and $Y_{i}=Y(A_{i})$ the realized reward. The actions are selected by the behavior policy $\pi_{0}$, where $\pi_{0}(a\,|\,x):=\mathbb{P}(A_{i}=a\,|\,X=x)$ is the propensity score, for any $a\in[M],x\in\mathcal{X}$. We make the following assumptions for $\pi_{0}$ and $P$.
Assumption 2.1.
@@ 2 Preliminaries
The above assumptions are standard in the literature (see e.g., Athey & Wager, 2021; Zhou et al., 2023; Si et al., 2023; Kallus et al., 2022). In particular, the unconfoundedness assumption guarantees identifiability, and the overlap assumption ensures sufficient exploration when collecting the training dataset. The bounded reward support is assumed for the ease of exposition, and can be relaxed to the sub-Gaussian reward straightforwardly.
@@ 2.1 The KL-distributionally Robust Formulation
We define an uncertainty set of neighboring distributions around $P$, whose conditional outcome distribution is bounded in KL divergence from $P$. Given a radius $\delta>0$, the uncertainty set of the conditional distribution is defined as $\mathcal{P}(P_{Y\,|\,X},\delta):=\big{\{}Q_{Y\,|\,X}:D_{\text{KL}}(Q_{Y\,|\,X}\,\|\,P_{Y\,|\,X})\leq\delta\big{\}}$, where $P_{Y\,|\,X}$ and $Q_{Y\,|\,X}$ refers to the distribution of $(Y(1),\dots,Y(M))\,|\,X$ under $P$ and $Q$ respectively. The distributionally robust policy value for any policy $\pi$ at level $\delta$ is defined as
@@
Under this formulation, our goal is to learn a “robust” policy with a high value of $\mathcal{V}_{\delta}(\pi)$ using a dataset drawn from $P$. The task here is two-fold: we need to (i) estimate the policy value $\mathcal{V}_{\delta}(\pi)$ for a given policy $\pi$, and (ii) find a near-optimal robust policy $\widehat{\pi}\in\Pi$ whose policy value is close to the optimal policy $\pi^{*}_{\delta}$. Here, the performance of a learned policy $\widehat{\pi}$ is measured by the sub-optimality gap (regret):